OpenStudy (anonymous):

1, 1, 2, 3, 5, 8, next 3 numbers?

ganeshie8 (ganeshie8):

fib...

OpenStudy (anonymous):

Its a fibonacci series.... try to add two previous nos. to get the new number

OpenStudy (anonymous):

Add the previous two numbers to get next number Next number is 5+8=13 Next number is 8+13=21 Last number is 13+21=34

OpenStudy (anonymous):

thanks! can someone help with this?? After a gas-filled balloon is released, it rises 90 feet by the end of the first minute. By the end of the second minute, the balloon rises 120 feet, and by the end of the third minute, it rises 150 feet. How many feet would the balloon rise in 8 minutes?

OpenStudy (anonymous):

welcome to open study please read the code of conduct http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

in each minutes balloon rises 30 feet so after 8 minutes rises 300 feet.. 90+30=120 120+30=150 150+30=180 180+30=210 ..... 270+30=300